### v2

# Generating Query Recommendations without Query Logs

Generating Query Recommendations without Query Logs
1. Results and Analysis
In this section, we present the results to answer our posed five research questions (Section LABEL:sec:research-questions).
RQ1: Can our proposed GQR system generate relevant and useful query recommendations compared with existing query recommendation systems?
In Tables 1 and 2 we report the clarity score and the NDCG@10 metrics, respectively, for the tested query recommendation systems, according to the Substitution evaluation protocol. For each system, collection, and metric, we report the minimum and maximum metric values across the six generated recommendations, as well as their average values and the standard deviations. According to the SCS metric, our GQR systems get the highest values with both the worst and the best performing recommendations on the Robust04 dataset, outperformed only by System 2’s best recommendation on the ClueWeb09B dataset, but by a small margin. However, on average, our proposed GQR (GPT-3) system outperforms all other methods, including the commercial systems. Moreover, the corresponding standard deviation values are the smallest, meaning that all the generated query recommendation obtain a similar, high SCS score. According to the NDCG@10 metric, our proposed GQR (GPT-3) system outperforms all other systems. With respect to the best competitor, namely System 2, our experiments show that GQR (GPT-3) has, on average, a +23% and a +27% improvement in NDCG@10 on Robust04 and ClueWeb09B, respectively. As with the SCS metric, the standard deviation of the NDCG@10 scores across the six generated recommendations is small, so the recommendations get a NDCG@10 score concentrated around the average value.
Model | Min | Max | Avg Std |
Robust04 | |||
System 1 (a) | 9.21 | 10.17 | 9.80 0.35 |
System 2 (b) | 9.35 | 10.56 | 9.82 0.39 |
Bhatia (bhatia2011query) (c) | 7.64 | 9.20 | 8.28 0.53 |
GQR (Bloom) (d) | 7.50 | 11.00 | 8.84 1.66 |
GQR (GPT-3) | 10.54abcd | 10.77c | 10.65 0.08abcd |
ClueWeb09B | |||
System 1 (a) | 9.80 | 10.80 | 10.37 0.32 |
System 2 (b) | 10.49 | 11.31 | 10.87 0.30 |
Bhatia (bhatia2011query) (c) | - | - | - |
GQR (Bloom) (d) | 7.68 | 11.20 | 9.84 1.19 |
GQR (GPT-3) | 10.94ad | 11.22 | 11.12 0.10abd |
Model | Min | Max | Avg Std |
Robust04 | |||
System 1 (a) | 0.2038 | 0.2638 | 0.2377 0.0211 |
System 2 (b) | 0.2546 | 0.3739 | 0.3102 0.0478 |
Bhatia (bhatia2011query) (c) | 0.2388 | 0.2640 | 0.2566 0.0108 |
GQR (Bloom) (d) | 0.1743 | 0.3655 | 0.2628 0.0714 |
GQR (GPT-3) | 0.3727 | 0.3947 | 0.3842 0.0092 |
ClueWeb09B | |||
System 1 (a) | 0.0940 | 0.1129 | 0.1056 0.0082 |
System 2 (b) | 0.0946 | 0.1294 | 0.1108 0.0164 |
Bhatia (bhatia2011query) (c) | - | - | - |
GQR (Bloom) (d) | 0.0468 | 0.1144 | 0.0802 0.0268 |
GQR (GPT-3) | 0.1280 | 0.1659 | 0.1403 0.0146 |
In Tables 3 and 4 we report the clarity score and the NDCG@10 metrics, respectively, for the tested query recommendation systems, according to the Concat evaluation protocol. This protocol aims to assess the performance improvements that a user query may get when the generated query recommendations are used as additional information sources together with the query. In doing so, the query recommendations are expected to better specify the information need of the user expressed as a text query. In doing so, it is important to take into account the number of query recommendations concatenated with the user query, so we report the metric scores for increasing number of recommendation concatenated, referreto to as rank. The average clarity score of the original queries is 12.03, respectively 12.50, on the Robust04, respectively ClueWeb09B, dataset. In both cases, the concatenation of one or more query recommendation increases the SCS value, for all tested systems, as expected. In fact, according to cronen2002predicting, the clarity score of a query increases if we add terms that reduce the query ambiguity. Since we are enriching the user queries with recommendations, these increases confirm that the generated recommendations do not increase the ambiguity of the queries. Among the tested systems, our proposed GQR (GPT-3) outperforms all other systems, across all ranks and datasets. Similar results are obtained w.r.t. NDCG@10. The NDCG@10 of the original queries is is 0.4341, respectively 0.1951, on the Robust04, respectively ClueWeb09B, dataset. In this case, more recommendations we concatenate to the query, the higher the metric value. With six recommendations, the effectiveness performance increases by on the Robust04 dataset and by on the ClueWeb09 dataset.
Concerning RQ1, we can conclude that our proposed GQR (GPT-3) systems is able to generate query recommendations that are, on average, less ambiguous than other system, and better or on par with commercial competitors. Moreover, the generated query recommendations can be used as additional information sources to encode the user information needs, with a resulting NDCG@10 improvement up to .
Model | rank 1 | rank 2 | rank 3 | rank 4 | rank 5 | rank 6 |
Robust04 | ||||||
System 1 (a) | 16.29 | 20.00 | 23.32 | 26.55 | 29.18 | 31.43 |
System 2 (b) | 19.35 | 24.81 | 31.10 | 36.63 | 42.84 | 47.88 |
Bhatia (bhatia2011query) (c) | 18.21 | 23.89 | 29.52 | 34.49 | 39.56 | 44.28 |
GQR (Bloom) (d) | 19.13 | 25.41 | 31.08 | 36.12 | 40.48 | 44.20 |
GQR (GPT-3) | 20.20ac | 27.69abcd | 34.98abcd | 42.23abcd | 49.47abcd | 56.67abcd |
ClueWeb09B | ||||||
System 1 (a) | 19.20 | 24.89 | 30.54 | 36.03 | 41.23 | 46.24 |
System 2 (b) | 19.28 | 24.50 | 30.77 | 35.55 | 41.55 | 46.70 |
Bhatia (bhatia2011query) (c) | - | - | - | - | - | - |
GQR (Bloom) (d) | 18.49 | 25.20 | 31.15 | 37.07 | 42.28 | 46.72 |
GQR (GPT-3) | 20.41d | 27.73abd | 35.01abd | 42.27abd | 49.30abd | 56.25abd |
Model | rank 1 | rank 2 | rank 3 | rank 4 | rank 5 | rank 6 |
Robust04 | ||||||
System 1 (a) | 0.4151 | 0.4270 | 0.4320 | 0.4373 | 0.4450 | 0.4417 |
System 2 (b) | 0.4236 | 0.4292 | 0.4361 | 0.4386 | 0.4433 | 0.4373 |
Bhatia (bhatia2011query) (c) | 0.4044 | 0.4012 | 0.3956 | 0.3917 | 0.3900 | 0.3922 |
GQR (Bloom) (d) | 0.4252 | 0.4254 | 0.4149 | 0.4136 | 0.4172 | 0.4203 |
GQR (GPT-3) | 0.4316 | 0.4429 | 0.4507 | 0.4543 | 0.4597 | 0.4596 |
ClueWeb09B | ||||||
System 1 (a) | 0.1598 | 0.1627 | 0.1832 | 0.1797 | 0.1912 | 0.1954 |
System 2 (b) | 0.1591 | 0.1429 | 0.1506 | 0.1528 | 0.1562 | 0.1645 |
Bhatia (bhatia2011query) (c) | - | - | - | - | - | - |
GQR (Bloom) (d) | 0.1586 | 0.1626 | 0.1576 | 0.1544 | 0.1443 | 0.1427 |
GQR (GPT-3) | 0.1960 | 0.1891 | 0.1919 | 0.1939 | 0.2010 | 0.2030 |
RQ2: Are the queries generated by our GQR system more engaging for users than those generated by other systems?
To answer this research question, we report the results of our user study in Table 5.
We exclude Bathia’s approach because it needs to index a collection of documents which is not provided with the AOL queries.
The user study shows a clear preference for our GQR (GPT-3) system with respect to the other two commercial systems. All groups report a preference above of tested queries for the GQR (GPT-3) system, with an overall preference of .
The second most preferred system is System 1 with of the preferences on average, and the last system is System 2, with of preferencess on average.
Hence, on RQ2, we can conclude that, according to our user study, the recommendations generated by our GQR (GPT-3) system are more engaging than the recommendations generated by the two commercial competitors from the user’s point of view.
It is worth noticing that AOL queries cover a wide variety of topics and since it is impossible to make assumptions about users’ prior knowledge and their personal preferences, for that reason, subjectivity plays a crucial role.
Annotator ID System 1 System 2 GQR (GPT-3) Group 1 Ann 1 20.31% 15.62% 64.06% Ann 2 32.81% 15.63% 51.56% Ann 3 31.25% 18.75% 50.00% Ann 4 26.98% 11.11% 61.90% Avg Std 27.84% 4.84% 15.28% 2.72% 56.88% 6.17% Group 2 Ann 5 20.31% 23.44% 56.25% Ann 6 39.06% 14.06% 46.88% Ann 7 25.00% 12.50% 62.50% Ann 8 31.25% 17.19% 51.56% Avg Std 28.91% 7.03% 16.80% 4.19% 54.30% 5.78% Group 3 Ann 9 17.19% 18.76% 64.06% Ann 10 26.56% 9.38% 64.06% Ann 11 18.75% 14.06% 67.19% Ann 12 25.00% 4.69% 70.63% Avg Std 21.88% 3.98% 11.72% 5.24% 66.41% 2.71% Overall 26.21% 6.26% 14.60% 4.69% 59.19% 7.33%
RQ3: Does our GQR system generate recommendations for long tail, i.e., rare, queries?
During our experiments on Robust04 and ClueWeb09B, we observed that some systems fail to produce query recommendations for some queries, while our GQR (GPT-3) system is always successful in generating at least 1 suggestion for each query.
To further investigate on this topic, we extracted 200 rare queries from the tail distribution of the AOL query log, thus considering the queries with a frequency equal to 1-2 in the query log.
In Table 6 we report, in percentage, the number of times our tested systems have been able to generate at least one and all six recommendations for the queries Robust04, ClueWeb09B and the 200 AOL tail queries datasets. We also report the average number of query recommendations each system successfully generated for each dataset.
For the Robust04 and ClueWbe09B datasets, all competitors are able to generate more than 5 recommendations on average. On average, all systems generate at least one recommendation for more than of the Robust04 queries, and for more than on ClueWeb09B. System 1 and System 2 are not able to generate at least one recommendation and of the times on AOL tail queries, respectively. Our GQR (Bloom) system generates at least one query recommendation of the times, but it is the worst performing system when generating all six recommendations. However, our GQR (GPT-3) system always succeeds in generating six recommendations on all tested datasets. So, concerning RQ3, we can conclude that our GQR (GPT-3) system always produces recommendations for all ranks.
Model | One suggestion | Six suggestions | Average |
Robust04 | |||
System 1 | 92% | 87% | 5.39 |
System 2 | 98% | 86% | 5.54 |
Bhatia | 95% | 89% | 5.51 |
Bloom | 99% | 55% | 4.69 |
GPT-3 | 100% | 100% | 6 |
ClueWeb09B | |||
System 1 | 98% | 92% | 5.75 |
System 2 | 99% | 95% | 5.85 |
Bloom | 99% | 63% | 5.1 |
GPT-3 | 100% | 100% | 6 |
AOL long tail queries | |||
System 1 | 83% | 71% | 4.61 |
System 2 | 91% | 86% | 5.37 |
Bloom | 99% | 66% | 5.15 |
GPT-3 | 100% | 100% | 6 |
RQ4: What is the impact of the number of examples in the prompt context on the performance of our GQR system?
To assess this RQ we conduct a prompt study on Robust04, in which we compare prompts with a different number of examples in it.
This prompt study analyzes the impact of the number of examples in the prompt on the performance in terms of SCS and NDCG@10
Our results, reported in Table 7, shows that the model with a single example in the prompt achieves comparable performances with the prompt with up to 10 examples.
Therefore we can conclude that the number of examples in a prompt is not relevant and that GQR still works using fewer examples.
Num. examples | SCS STD | NDCG@10 STD |
1 | 10.26 0.102 | 0.3888 0.007 |
2 | 10.50 0.134 | 0.3816 0.007 |
5 | 10.47 0.077 | 0.3785 0.008 |
10 | 10.65 0.08 | 0.3842 0.009 |
RQ5: What is the impact of the specific prompt context on the performance of our GQR system?
To understand if the performances of GQR are conditioned by the specific examples provided in the prompt, we conduct another prompt study on Robust04.
This study tests ten prompts with two examples, each of which is unique. For instance, a prompt will be composed with examples related to a celebrity and an airline company and another one with a consumer product and a political event.
Table 8 shows that the various prompts achieve very similar effectiveness performance and they are not statistically different. Hence we can conclude that the specific query-list of recommendations prompt is not conditioning the GQR (GPT-3) generation process.
Alternative Prompt | SCS STD | NDCG@10 STD |
1 (ours) | 10.48 0.08 | 0.3816 0.007 |
2 | 10.47 0.11 | 0.3838 0.009 |
3 | 10.83 0.09 | 0.3908 0.010 |
4 | 10.89 0.17 | 0.3895 0.011 |
5 | 10.55 0.05 | 0.3919 0.005 |
6 | 10.61 0.02 | 0.393 0.004 |
7 | 10.87 0.08 | 0.3855 0.014 |
8 | 10.73 0.08 | 0.3929 0.009 |
9 | 10.69 0.10 | 0.3798 0.004 |
10 | 10.39 0.11 | 0.3896 0.013 |